OpenStudy (anonymous):

The typical American generates about 3.8 pounds of solid trash per day with a standard deviation of 1.5 pounds. Assuming that the histogram for trash produced per day is approx. bell shaped, between what two values will 95% of the observation fall? Thanks

OpenStudy (anonymous):

From the empirical rule, about 95% of the data will lie between: \[\mu -2\sigma\]and\[\mu+2\sigma\]So your lower bound will be approximately,\[3.8 - 2 \times 1.5 = 0.8\]and\[3.8+2 \times 1.5=6.8\]Hence, 95% of the observations would fall between 0.8 lb and 6.8 lb.